A cylinder measuring 6.1 cm wide and 7.3 cm high is filled with gas and the piston pushed down with a steady force piston cylinder gas Calculate the force on the piston if the pressure in the gas is measured to be 314. kPa, Write your answer in units of kilonewtons. Be sure your answer has the correct number of significant digits. kN ×1
Answer #1

1)

P = 314 kPa

= 314*10^3 Pa

= 3.14*10^5 Pa

Area of upper part = pi*(diamter/2)^2

= 3.14*(6.1/2 cm)^2

= 3.14*(3.05 cm)^2

= 29.2 cm^2

= 2.92*10^-3 m^2

Now use:

P = F/A

3.14*10^5 Pa = F / (2.92*10^-3 m^2)

F = 917 N

F = 0.917 KN

Answer: 0.92 KN

2)

Molar mass of CO,

MM = 1*MM(C) + 1*MM(O)

= 1*12.01 + 1*16.0

= 28.01 g/mol

mass(CO)= 19 g

use:

number of mol of CO,

n = mass of CO/molar mass of CO

=(19 g)/(28.01 g/mol)

= 0.6783 mol

Given:

V = 5.0 L

n = 0.6783 mol

T = 30.0 oC

= (30.0+273) K

= 303 K

use:

P * V = n*R*T

P * 5 L = 0.6783 mol* 0.08206 atm.L/mol.K * 303 K

P = 3.3731 atm

Answer: 3.4 atm